clear fractions, solve and check for problem 2/3+1/4t=2
Thank you

Answer by ewatrrr(24785)

 
Hi
Solve by isolating 't' on one side of the equation
2/3 + (1/4)t = 2
2/3 + (1/4)t = 6/3
(1/4)t = 4/3
t = 16/3 or 5 1/3
CHECKING our Answer***
2/3 + (1/4)*16/3 = 2/3 + 4/3 = 6/3 = 2



ewatrrr@aol.com
private/pay tutoring Inquiries only, please.